### v1

# Take Fake as Real: Realistic-like Robust Black-box Adversarial Attack to Evade AIGC Detection

Abstract
The security of AI-generated content (AIGC) detection based on GANs and diffusion models is closely related to the credibility of multimedia content. Malicious adversarial attacks can evade these developing AIGC detection. However, most existing adversarial attacks focus only on GAN-generated facial images detection, struggle to be effective on multi-class natural images and diffusion-based detectors, and exhibit poor invisibility. To fill this gap, we first conduct an in-depth analysis of the vulnerability of AIGC detectors and discover the feature that detectors vary in vulnerability to different post-processing. Then, considering the uncertainty of detectors in real-world scenarios, and based on the discovery, we propose a Realistic-like Robust Black-box Adversarial attack (R2BA) with post-processing fusion optimization. Unlike typical perturbations, R2BA uses real-world post-processing, i.e., Gaussian blur, JPEG compression, Gaussian noise and light spot to generate adversarial examples. Specifically, we use a stochastic particle swarm algorithm with inertia decay to optimize post-processing fusion intensity and explore the detector’s decision boundary. Guided by the detector’s fake probability, R2BA enhances/weakens the detector-vulnerable/detector-robust post-processing intensity to strike a balance between adversariality and invisibility. Extensive experiments on popular/commercial AIGC detectors and datasets demonstrate that R2BA exhibits impressive anti-detection performance, excellent invisibility, and strong robustness in GAN-based and diffusion-based cases. Compared to state-of-the-art white-box and black-box attacks, R2BA shows significant improvements of 15% and 21% in anti-detection performance under the original and robust scenario respectively, offering valuable insights for the security of AIGC detection in real-world applications.
Index Terms:
Adversarial attack, AIGC detection, Post-processing, Particle swarm optimization
I Introduction
With the rapid development of GAN- and diffusion-based generation technologies[1], AI-generated content (AIGC) detection has become an increasingly popular research topic[2],[3],[4]. The security of AIGC detection is closely linked to the credibility of multimedia content[5]. If the security of AIGC detection technology is not guaranteed, multimedia content will face a serious trust crisis. However, AIGC detection technologies are still in an immature stage: Many studies[6],[7],[8],[9] have shown that GAN-based image forgery detectors are vulnerable to malicious adversarial attacks, resulting in detection failures. On the other hand, the widespread use of diffusion models makes security research on diffusion-based detection increasingly urgent. However, the lack of robustness research on this detection method[2],[4],[10],[11],[12], combined with the fact that the generative mechanisms of diffusion models are completely different from those of GANs[13],[14], leaves the security of diffusion-based AIGC detection uncertain. Therefore, to enhance the security of AIGC, it is crucial to focus on the robustness of its detection against adversarial attacks, particularly those diffusion-based methods.
Figure 1: Principle of our method. We use Gaussian blur, JPEG compression, Gaussian noise and light spot to process AI generated images. The post-processed image serves as an adversarial example that leads the detector to recognize the AI-generated image as a real/non-AI-generated image.
The existing adversarial attacks against AIGC detectors primarily focus on GAN-based detectors. Some studies[6],[7],[8],[15] employ gradient-based attacks to generate perturbations, but these perturbations are always recognized as fake by well-trained detectors. Additionally, these studies[6],[7],[8],[9],[15] perform well on Deepfake facial datasets but poorly on multi-class natural image datasets, which restricts their broader application. Furthermore, the invisibility of existing adversarial examples[9],[16] is often overlooked, leading to severe image distortions that limit their practical use. Therefore, designing realistic-like adversarial attacks applicable to multi-class natural images is of significant practical value for deploying AIGC detectors in real-world scenarios.
Previous researches[9],[17],[18],[19],[20] demonstrate that adversarial examples can be generated by employing natural degradation as an attack method. These methods generate adversarial examples by simulating the degradation of image quality in the real world, which is more meaningful for studying the robustness of the AIGC detectors, as these post-processing techniques, like JPEG compression and Gaussian blur, are common on social networks. In other words, if a detector classifies such post-processed AIGC images as real, it suggests that the detector will face similar performance degradation issues in real-world applications. Therefore, employing post-processing techniques to generate adversarial samples is both feasible and of significant practical value.
In order to design a realistic-like adversarial attack applicable to natural images, as shown in Fig. 1, we select Gaussian blur, JPEG compression, Gaussian noise, and light spot as attacks to generate adversarial examples. These 4 techniques have been proven to attack the frequency, spatial, and statistical domains of detectors, respectively[9],[17],[19],[20]. And the pre-experiment in Fig. 2 reveal the discovery that detectors vary in vulnerability to different post-processing. Based on the discovery, and considering the uncertainty of detectors in real-world scenarios, we propose R2BA, a realistic-like robust black-box adversarial attack based on post-processing fusion optimization. During the optimization process, R2BA first randomly generates multiple post-processed adversarial examples to explore the detector’s complex decision boundary. Then, guided by the fake probability output of the AIGC detector, R2BA continuously updates the post-processing intensity of these adversarial examples, driving them toward and eventually across the decision boundary of the detector, thereby evading AIGC detection.
R2BA solves 3 problems existed in previous work:
Applicability to AIGC detectors. Based on the discovery that detectors vary in vulnerability to different post-processing, we integrate 4 widely-used post-processing techniques to launch a comprehensive attack, ensuring all detectors are successfully attacked by R2BA, resulting in significant performance degradation.
Balance between invisibility and adversariality. Guided by the detector’s fake probability, R2BA enhances the detector-vulnerable post-processing intensity to improve adversariality, and weakens the detector-robust post-processing intensity to minimize unnecessary modifications, making the post-processed AIGC images appear highly invisible and realistic from a human visual perspective.
Multi-class natural images effectiveness. R2BA is not a class-wise method and does not set any constraints on data distribution. To address the complex decision boundary issues introduced by multi-class natural images, R2BA randomly initialize multiple post-processed images to expand the search space. This approach allows us to efficiently locate the decision boundaries that are most vulnerable to the fusion post-processing, thus evading detection.
Compared to existing adversarial attacks against AIGC detectors, R2BA generates adversarial examples with real-world significance. It demonstrates powerful anti-detection performance on AIGC detectors for natural images, not limited to facial images or GAN-based detectors, thereby filling the gap in adversarial attacks against both GAN-based and diffusion-based AIGC detectors.
Our contributions can be summarized as follows:
  • We propose R2BA, a novel realistic-like robust black-box adversarial attack, which is the first method with significant practical value, capable of evading AIGC detectors on natural images. It demonstrates applicability to black-box detectors, with impressive anti-detection performance, excellent image quality, and strong robustness.
  • We propose a post-processing fusion optimization to achieve a trade-off between adversariality and invisibility. Through a stochastic particle swarm optimization algorithm with inertial decay, and oriented to the detector vulnerability to different post-processing, this optimization realizes the dynamic adjustment of the fusion intensity.
  • We conduct extensive experiments on 3 popular open-source AIGC image datasets, validating the effective anti-detection performance of R2BA in evading state-of-the-art GAN-based and diffusion-based AIGC detectors, as well as commercial AIGC detectors, in the original, robust and real-world scenarios.
The remainder of this paper is organized as follows. Section II reviews related work. Section III presents the specific implementation of the proposed method. The experiments and analyses are described in Section IV. Finally, we provide a brief conclusion in Section V.
II Related Work
II-A AI-generated Image Forgery Detection
AI-generated image forgery detection is actually a Fake-Real binary classification problem. Here, Fake refers to AIGC, which means the image is generated by AI, specifically through techniques like GANs or diffusion models. While Real refers to UGC, meaning that the image is a natural image or user-generated. The challenge of AIGC detection lies in extracting subtle forgery signals from AI-generated images and making accurate distinctions. According to different generation technologies, AIGC detection techniques can be divided into GAN based and diffusion based methods.
GAN-based forgery detection. Image generation based on GAN architecture[21],[22],[23] consists of generative and discriminative models. The goal of this architecture is to train the generative model to exactly fit the real data distribution, so that the discriminative model is difficult to distinguish. To prevent this technique from being misused, several GAN-based forgery detection techniques[10],[11],[24],[25],[26] have been proposed. These works recognize AI-generated images by extracting specific features and use data-driven approaches to guide the model in learning the differences in features across the spatial, frequency, and statistical domains between AI-generated and natural images.
Diffusion-based forgery detection. Diffusion-based image generation[27],[28],[29] has gained considerable attention due to its powerful generative capabilities. Unlike the GAN architecture, diffusion models completely revolutionize the traditional generative process. These models use high-dimensional latent variables to capture more details of the data, resulting in higher-quality images. It is evident that the powerful generative ability of diffusion models enhances the robustness of diffusion-based forgery detection techniques[2],[4],[10],[11]. These techniques focus on capturing subtle forgery signals in diffusion-generated images through methods such as reconstruction or contrast learning, aiming to identify small differences that may be exist in the image, and thus accurately distinguishing between real and fake content.
II-B Adversarial Attacks for Anti-detection
In order to explore the robustness of detectors, many adversarial attacks have been used to generate adversarial examples for anti-detection. These works aim to generate adversarial examples for AI-generated images, which cause the detector to misclassify the fake/AI-generated images as real images. However, current anti-detection adversarial attacks[6],[7],[9] are mainly performed on the Deepfake face dataset, which limits their broader applications and often results in severe performance degradation when against robust AIGC detectors on the multi-class AI-generated natural image datasets. The following are some popular adversarial attack methods.
White-box attack. Classical white-box attack methods[6],[7],[8] rely on obtaining gradient information from the target model to generate adversarial examples. But the meaninglessness and randomness of these perturbations may make such algorithms fail to attack in more complex tasks or scenarios. Some studies[9],[15] aim to evade GAN-based detection through statistical feature consistency or inconspicuous attribute variation in facial images. However, these methods often fail because it is difficult to determine the features that the detector truly focuses on, and the attributes of natural images are highly diverse, with significant differences across different classes of images.
Black-box attack. Black-box adversarial attacks do not require access to the inner workings of the model, and efficient attacks can be achieved by only obtaining the model output. There are three main types of black-box attacks: Algorithms based on gradient estimation[7] and agent models[30] are limited due to their reliance on fitting the target. When applied in practice, query-based methods might be the best solution. However, when facing the unpredictable decision boundaries caused by multi-class natural images[31], a single adversarial example requires significant effort to locate the decision boundary. Some methods based on random pixel updates[16],[32] are effective in anti-detection, but tend to cause significant perturbations, which allows these adversarial examples to be directly classified as forged images by the human eye, even without a detector. Another method involving attribute editing[15], although optimizing multiple adversarial examples simultaneously, still face the challenge of selecting the specific attribute across the diverse natural images and requires appropriate attribute editing tools.
Based on the above analysis, existing adversarial attacks for anti-detection have limitations in image classes (they only work on facial datasets) and detectors (primarily GAN-based detection). Additionally, they are often difficult to strike a balance between the anti-detection effectiveness of adversarial examples and image quality. Moreover, in practical applications, real-world platform image processing may erase the adversarial perturbations, leading to the failure of anti-detection. To address the issues in existing works, we conduct comprehensive experiments on the security of GAN-based and diffusion-based AIGC detectors using natural image datasets with multiple classes, and we thoroughly analyze and compare various anti-detection performance metrics of our method with existing methods in original, robust and real-world scenarios.
Figure 2: Average performance degradation. The values represent the attack success rate of different post-processing against AIGC detectors. Darker colors indicate greater vulnerability of detectors. Attacks used include Gaussian blur, JPEG compression, Gaussian noise, light spot, and sequential fusion attacks.
III Proposed Method
In this section, we first explain the motivation for R2BA in subsection III-A. Next, we use a visualization in subsection III-B to illustrate how R2BA generates adversarial examples by optimizing post-processing intensity. Finally, we provide a detailed description of the R2BA implementation in subsection III-C, including specific formulas and the algorithm.
III-A Motivation
Traditional adversarial examples[6],[7],[16],[32] generated by gradient-based attacks or random pixel modifications attacks, are both meaningless and ineffective because their poor invisibility makes them difficult to apply in real-world scenarios. They can be easily identified as fake by the human eye, without the need for a detector. In contrast, using post-processing techniques commonly employed in real-life scenarios to create adversarial examples makes them look realistic and match real natural image patterns. Furthermore, if a detector classifies such post-processed AIGC images as real, this suggests that the detector will face similar performance degradation in real-world applications.
Inspired by previous works[9],[18],[20], we select Gaussian blur, JPEG compression, Gaussian noise and light spot as attacks, which have been proven to mislead detectors through frequency, spatial, and statistical domains respectively. Pre-experiment we conduct in Fig. 2 further confirm that these post-processing methods lead to varying degrees of performance degradation of detectors. We use Gaussian blur with a radius of $5$ and deviation of $1.5$, JPEG compression with an intensity of $80$, Gaussian noise with standard deviation of $7/255$, and light spot with an intensity of $1.8$ and random radius. The datasets[4],[24] and detectors[4],[10],[11],[24],[25] used in Fig. 2 are described in Section IV.
The results in Fig. 2 demonstrate that detectors vary in vulnerability to different post-processing. For example, on the CNNSpot[24] dataset, the CNNSpot[24] detector is highly vulnerable to Gaussian noise but more robust to JPEG compression (GN=0.12, JG=0.00), while the FatFormer[11] detector exhibits the opposite behavior (GN=0.01, JG=0.23). Interestingly, the effectiveness of post-processing fusion attacks does not always exceed that of individual post-processing on different detectors. On the DRCT[4] dataset, the attack performance of Gaussian blur + JPEG compression on the FatFormer[11] is better than that of Gaussian blur alone (GB+JG=0.87, GB=0.18), whereas the opposite (GB+JG=0.02, GB=0.12) is true for the UnivFD[10]. Therefore, optimizing the intensity of post-processing fusion is challenging, but necessary.
Therefore, considering the uncertainty of detectors in real-world scenarios, we adopt a black-box attack to generate adversarial examples. To attack different types of detectors, we fuse these four post-processing and define the attack order as step $1$-$4$ in Fig. 1. To solve the challenge of optimizing the fusion intensity, we use the detector’s fake probability as guidance and apply stochastic particle swarm optimization (PSO)[33] with inertia decay to optimize the fusion intensity. Compared to other optimization algorithms, such as evolutionary algorithms and simulated annealing, stochastic PSO with inertia decay outperforms them in its ability to effectively perform high-dimensional global searches[34]. This method can quickly locate the global decision boundary and simultaneously optimize fusion intensity to balance invisibility and adversariality. All things considered, we propose a post-processing fusion optimization-based, realistic-like robust black-box adversarial attack method, R2BA, to study the security of AIGC detectors in real-world applications.
III-B Visualisation
Figure 3: The pipeline of R2BA. It contains 4 modules and a main optimization process (Initialization-Optimization iteration-Select). R2BA uses N randomly generated post-processing parameters as particle positions, optimizing their velocities and positions to balance the invisibility and adversariality. A successful attack occurs when a particle crosses the detector’s decision boundary, and R2BA outputs the adversarial example with the highest SSIM value among the successful ones. The notations’ definitions here are the same as in Table I.
Component description. In Fig 3, we illustrate the pipeline of R2BA. It mainly consists of four modules: the post-processing module, fitness calculation, velocity/position update, and the swarm optimization module. The decision boundary here refers to the fake threshold ($0.5$) of the detector, which typically classifies an image with a fake probability greater than $0.5$ as fake and less than $0.5$ as real.
1) Post-processing Module: The red circle represents an AIGC image and the blue represent its post-processed images (adversarial particles). The particle positions represent the post-processing fusion intensity. We can balance invisibility and adversariality by optimizing the particle positions.
2) Fitness Calculation: The post-processed image is fed into the detector to obtain its fake probability, which serves as its particle’s fitness. A smaller fitness value indicates a more effective adversarial example. The individual best of each particle has the smallest fitness in the optimization process, i.e., the lowest fake probability achieved so far. The global best is the particle with the smallest fitness among all individual bests. The global best particle is most likely to cross the decision boundary, as it has the lowest fake probability.
3) Velocity / Position Update: Velocity $V_{i}^{k+1}$ updates are influenced by the particle’s individual best $p_{best}$, global best $P_{best}$, and inertia $w^{k}$. While $p_{i}^{k+1}$ position updates depend on the current $p_{i}^{k}$ position and velocity $V_{i}^{k+1}$. Obviously, influenced by the individual best and global best, the particles will move towards positions with a lower fake probability, continuously approaching the decision boundary.
4) Swarm Optimization Module: To optimize the particles’ positions (post-processing intensity) and reduce their fake probability, all particles should undergo fitness calculation as well as velocity and position updates.
Optimization Iteration. In the first iteration, to cover as much of the example domain space as possible, R2BA randomly assigns $N$ post-processing fusion intensity values as the initial positions of $N$ particles. Stochastic generation helps quickly locate the detector’s most vulnerable decision boundary, enabling efficient global search and addressing the challenges posed by diverse decision boundaries in multi-class natural images. In subsequent iterations, take the $k$-th iteration as an example, guided by the fitness, the particle velocity gradually points towards the decision boundary, with the swarm becoming more concentrated. In the final (or successful) iteration, some particles cross the boundary, causing the detector to misclassify them as real images. R2BA then output the adversarial example with the highest SSIM among all the particles with successful attacks.
Special designs. It is clear that the optimization process largely depends on whether the particles can locate the detector’s decision boundary to execute the attack. To improve search capacity, we introduce inertia decay and random velocity updates to expand the search space and avoid local optima.
The above description outlines a typical iterative process. However, in practice, AIGC images may be located near the decision boundary, causing the attack to succeed in the first iteration. If the images are far from the decision boundary, the attack may fail to cross it within the range defined by the post-processing intensity settings. Experiments show that the former case is more common, while the latter is relatively rare.
III-C Implementation
This subsection provides the specific implementation of R2BA, with the pseudocode shown in Algorithm 1. The main notations are summarized in Table I.
TABLE I: Definitions of Main Notations
Notations Definition
$D$()($\cdot$)( ⋅ ) Victim detector with output as forgery probability
$N$ The number of particles
$K$ The max number of iterations
$P^{k}$ The $k$-th iteration of the particle swarm
$w^{k}$ The inertia of the $k$-th iteration
$[W_{\text{min}},W_{\text{max}}]$ The range of inertia
$C_{\text{p}}$, $C_{\text{total}}$ Individual confidence and particle swarm confidence
$Pos$()($\cdot$)( ⋅ ) The position of a particle
$Pos_{\text{best}}$()($\cdot$)( ⋅ ) Optimal position of one particle or particles
$f_{i}^{k}$ Fitness of the $i$-st particle of the $k$-st iteration
$f_{i,\text{best}}$, $f_{\text{best}}$ Optimal fitness of one particle / particles
$\alpha,\beta$ The radius and variance of Gaussian blur
$\phi$ The compressive strength of JPEG compression
$\sigma$ The variance of Gaussian noise
$(\lambda_{x},\lambda_{y})$ The coordinates of light spot center
$\theta,\gamma$ The center brightness and radius of light spot
Particle swarm initialization. First, the algorithm initializes the global parameters $N$, $K$, $W_{\text{min}}$, $W_{\text{max}}$, $C_{\text{p}}$, $C_{\text{total}}$. Then, to avoid weak intensity that leads to poor adversariality, or excessive intensity that causes image distortion, set the range for the post-processing intensity. We formulate the $k$-th particle swarm as $P^{k}=(p_{1}^{k},p_{2}^{k},\dots,p_{N}^{k})$.
Each particle in the particle swarm, such as $p_{i}^{k}$, the $i$-th particle of $k$-th iteration, has four attributes:
1) The position $Pos(p_{i}^{k})$: That is the post-processing fusion intensity of the particle, which is formulated as Eq. 1.
$Pos(p_{i}^{k})=(\alpha,\beta,\phi,\sigma,\lambda_{x},\lambda_{y},\theta,\gamma),$ (1)
where $k=1$ and these 8 parameters symbolize the intensity of the four post-processing, randomly generated by Eq. 2. In Eq. 2, $\tau\in\{\alpha,\beta,\phi,\sigma,\lambda_{x},\lambda_{y},\theta,\gamma\}$, $\tau_{\text{min}}$ and $\tau_{\text{max}}$ is the lower bound and upper bound on the corresponding parameter.
$\tau=random(\min=\tau_{\text{min}},\max=\tau_{\text{max}})$ (2)
Algorithm 1 R2BA optimization process.
0:  Original fake image $\mathbf{X}$
0:  Adversarial example $\mathbf{X}_{\text{adv}}$
1:  Initialize global parameters $N$, $K$
2:  Set $W_{\text{min}}$, $W_{\text{max}}$, $C_{\text{p}}$, $C_{\text{total}}$
3:  Set the ranges of $\alpha,\beta,\phi,\sigma,\lambda_{x},\lambda_{y},\theta,\gamma$
4:  Initialize particles by Eq.1 to Eq.4
5:  for $i=1,2,\dots,N$ do
6:     Calculate fitness $f_{i}^{1}$ of $p_{i}^{1}$ with $\mathbf{X}$ by Eq.5
7:     Set $f_{i,\text{best}}$ and $Pos_{\text{best}}(p_{i})$ by Eq.6 and Eq.7
8:  end for
9:  $f_{\text{best}}=\min_{\begin{subarray}{c}i=1,2,\dots,N\end{subarray}}\{f_{i,\text{best}}\}$
10:  set $Pos_{\text{best}}(P)$ corresponding to $f_{\text{best}}$
11:  Set current iteration $k=2$
12:  while $k\leq K\text{ and }f_{\text{best}}\geq 0.5$ do
13:     for $i=1,2,\dots,N$ do
14:        Update $w^{k}$, $v_{i}^{k}$, $Pos(p_{i}^{k})$ by Eq. 8 to Eq. 13
15:        Calculate fitness $f_{i}^{k}$ of $p_{i}^{k}$ with $\mathbf{X}$ by Eq.5
16:        if $f_{i}^{k}<f_{i,\text{best}}$ then
17:           $f_{i,\text{best}}=f_{i}^{k}$
18:           $Pos_{\text{best}}(p_{i})=Pos(p_{i}^{k})$
19:        end if
20:        if $f_{i,\text{best}}<f_{\text{best}}$ then
21:           $f_{\text{best}}=f_{i,\text{best}}$
22:           $Pos_{\text{best}}(P)=Pos_{\text{best}}(p_{i})$
23:        end if
24:     end for
25:  end while
26:  Select $Pos_{\text{adv}}(P)$ by Eq. 18 and Eq. 19
27:  $\mathbf{X}_{\text{adv}}$= $\mathbf{X}_{Pos_{\text{adv}}(P)}$
28:  return  $\mathbf{X}_{\text{adv}}$
2) The velocity $v_{i}^{k}$: The velocity of each particle will be initialized according to Eq. 3 and Eq. 4.
$v_{i}^{k}=(v_{i(1)}^{k},v_{i(2)}^{k},\dots,v_{i(M)}^{k}),M=8$ (3)
$v_{i(j)}^{1}=random(\min=-1,\max=1)$ (4)
where $v_{i}^{k}$ is the velocity of $i$-th particle of the $k$-th iteration, and $v_{i(j)}^{k}$ is the $j$-th latitude velocity of $v_{i}^{k}$. $M$ is the dimension of the problem. Since we have 8 parameters, $M=8$.
3) The fitness $f_{i}^{k}$: We calculate the fitness of each particle by Eq. 5. $X_{Pos(p_{i}^{k})}$ denotes the post-processed image of $X$, and $Pos(p_{i}^{k})$ is the post-processing intensity.
$f_{i}^{k}=D(X_{Pos(p_{i}^{k})})$ (5)
where $D(\cdot)$ is the victim detector with a fake probability output. In other words, a smaller fitness value indicates a lower fake probability, implying better adversariality.
4) The individual best $f_{i,\text{best}}$ and $Pos_{\text{best}}(p_{i})$: The individual best, which refers to the best position $Pos_{\text{best}}(p_{i})$ corresponding to its all-time best fitness $f_{i,\text{best}}$. In initialization, we set the individual best by Eq. 6 and Eq. 7. and define the particle with the smallest fitness as the global best $f_{\text{best}}$ and $Pos_{\text{best}}(P)$.
$f_{i,\text{best}}=f_{i}^{1},i=1,2,\dots,N$ (6)
$Pos_{\text{best}}(p_{i})=Pos(p_{i}^{1}),i=1,2,\dots,N$ (7)
Optimization iteration process. Now we have completed the initialization. There comes the optimization iteration. Before each iteration begins, we need to determine particle’s current position $Pos(p_{i}^{k})$ and its next velocity $V(p_{i}^{k+1})$, which are the key factors that influence the particle’s next position $Pos(p_{i}^{k+1})$. Here the former is known. Therefore, the focus is on how to update the velocity of the particles. The update of the particle’s velocity involves two aspects: inertia $w$ and the two best $Pos_{\text{best}}(p_{i})$ and $Pos_{\text{best}}(P)$. The following describes how to update the particle’s velocity and position respectively.
1) Velocity update
The inertia weight $w$ control the extent to which a particle depends on its previous velocity during updates. By gradually decreasing the inertia weight, R2BA can take different convergence modes during the optimization: Early in the optimization, larger inertia weight accelerates global exploration, enhances the ability to explore the entire example domain space, and reduces the risk of local optima. As optimization progresses and inertia decreases, particles can more precisely converge on the global best, reducing unnecessary position shifts and improving both convergence speed and accuracy, while also enhancing the ability to cross the decision boundary.
In R2BA, using linear decay allows the particles to gradually and smoothly towards the global best solution. The formula for inertia decay is given in Eq. 8.
$w^{k}=W_{\text{max}}-(W_{\text{max}}-W_{\text{min}})*\frac{k}{K}$ (8)
The particle’s velocity controls its movement. We want the particle to move towards the decision boundary (where the fake probability is 0.5) and even across it to execute a successful attack. Therefore, it is the most direct and effective method to guide the velocity update with the individual best and the global best, which are always updated to the position with the smallest fake probability. Therefore, we use the distance between $Pos(p_{i}^{k})$ and $Pos_{\text{best}}(p_{i})$/$Pos_{\text{best}}(P)$ as guidance to update the particle’s velocity, which in turn influences the particle’s position in the next iteration. Combined with the inertia $w$, the velocity is update by Eq. 9, Eq. 10 and Eq. 11.
$v_{i}^{k+1}=w^{k}\cdot v_{i}^{k}+C_{\text{p}}\cdot r_{\text{p}}\cdot V_{\text{p}}+C_{\text{{total}}}\cdot r_{\text{{total}}}\cdot V_{\text{{total}}}$ (9)
$V_{\text{p}}=Pos_{\text{best}}(p_{i})-Pos(p_{i}^{k})$ (10)
$V_{\text{{total}}}=Pos_{\text{best}}(P)-Pos(p_{i}^{k})$ (11)
$V_{\text{p}}$ and $V_{\text{{total}}}$ represent the influence of the individual best position and the global best position from $k$-th iteration on the particle’s velocity in $k+1$-th iteration, respectively.
2) Position update
Next, we update the particle’s position based on the updated velocity $v_{i}^{k+1}$, the previous position $Pos(p_{i}^{k})$ and a random modification term $\varepsilon_{i}$ by Eq. 12 and Eq. 13. The particle velocity incorporates both individual best and global best experiences, which are the key factors driving the particle towards the decision boundary.
$Pos(p_{i}^{k+1})=Pos(p_{i}^{k})+v_{i}^{k+1}+\varepsilon_{i}$ (12)
$\varepsilon_{i}=\begin{cases}random(),&\text{with prob }\epsilon,\quad r\sim U(0,1),\\ 0,&\text{with prob }1-\epsilon\end{cases}$ (13)
Here $\varepsilon_{i}$ is updated every iteration to increase the diversity of particles. And we set probability $\epsilon$ to 0.5, and $random()$ can output a decimal in the range $[0,1)$ for modification.
Since the updated positions may exceed the defined parameter range, we need to set constraints on the particles’ positions. The basic constraints are given in Eq. 14. For certain parameters, such as the radius and coordinates, we need to ensure that they are integers or odd numbers. These constraints are described in Eq. 15 and Eq. 16.
$\tau=\min(\max(\tau,\tau_{\text{min}}),\tau_{\text{max}})$ (14)
$\tau=Integer(\tau),\tau\in\{\alpha,\phi,\lambda_{x},\lambda_{y},\gamma\}$ (15)
$\tau=Odd(\tau),\tau\in\{\alpha\}$ (16)
The updates have now been completed. And next we follow the steps outlined in lines 15 to 23 of Algorithm 1 to optimize the individual best of particles and the global best of the particle swarm. Our optimization goal is to minimize the fake probability of all adversarial particles, as shown in Eq. 17.
[…]
Specifically, we use Eq. 5 to calculate the fake probability as the particle’s fitness, and compare them one by one to determine the individual best and the global best.
If the global best fitness is greater than 0.5 and the iteration has not reached the max iteration, the algorithm proceeds to the next optimization iteration. Otherwise, go to selection.
Selection. We use Eq. 18 to determine the final output $Pos_{\text{adv}}(P)$. When the global best fitness is larger than $0.5$, we let $Pos_{\text{adv}}(P)$ to be the particle with the smallest fitness. Otherwise, we select the particle with the largest SSIM[35] from the particles with successful attacks. The $ssim(\cdot)$[35] calculate the structural similarity between images, where the larger the value, the better the image quality.
$Pos_{\text{adv}}(P)=\begin{cases}Pos_{\text{best}}(P),&f_{\text{best}}\geq 0.5,\\ Pos_{\text{ssim}},&otherwise\end{cases}$ (18)
$Pos_{\text{ssim}}=Pos_{max\{ssim(\mathbf{X}_{Pos_{\text{best}}(p_{i})},\mathbf{X})|f_{i,\text{best}}<0.5\}}$ (19)
IV Experiments
TABLE II: GAN-based and Diffusion-based AIGC Datasets
Real Source Fake Type Models Data source
ImageNet[36] GAN ProGAN[22] CNNSpot[24]
ImageNet[36] Diffusion VQDM[28] GenImage[3]
MSCOCO[37] Diffusion LDM[27] DRCT[4]
IV-A Experimental Setups
Datasets. We conduct experiments on 3 datasets to evaluate our method. For GAN-based forgery detection, we use CNNSpot[24] as the GAN-based AIGC dataset. For diffusion-based forgery detection, we use GenImage[3] and DRCT[4] as the diffusion-based AIGC datasets. All 3 datasets are generated using publicly available forgery models. Table II shows the statistics of the AIGC datasets we used.
Victim detectors. We use CNNSpot[24], ResNetND[25], UnivFD[10], and FatFormer[11] for GAN-based detection, and FatFormer[11], UnivFD pretrained by DRCT[4], and ConvB[12] pretrained by DRCT[4] for diffusion-based detection. Here the FatFormer keeps the same on all datasets. Though UnivFD and UnivFD(DRCT) share the same model architecture, the specific models employed are different. We show these detectors’ accuracy in Table III and Table IV.
TABLE III: The accuracy of GAN-based AIGC detectors
Datasets Detectors
CNNSpot[24] ResNetND[25] UnivFD[10] FatFormer[11]
CNNSpot[24] 0.993 0.947 0.997 0.998
TABLE IV: The accuracy of Diffusion-based AIGC detectors
Datasets Detectors
FatFormer[11] UnivFD(DRCT)[4] ConvB(DRCT)[4]
GenImage[3] 0.810 0.891 0.769
DRCT[4] 0.963 0.967 0.999
TABLE V: The attack success rate and image quality results on GAN-based AIGC detectors
Dataset Attack CNNspot[24] ResNetND[25] UnivFD[10] FatFormer[11] Average
ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑
CNNSpot[24] FGSM[38] 0.565 42.367 0.687 0.590 42.453 0.688 0.015 40.071 0.704 0.131 40.364 0.703 0.325 41.314 0.696
MIFGSM[39] 0.575 42.137 0.688 0.570 42.596 0.688 0.020 40.624 0.703 0.090 40.339 0.703 0.314 41.424 0.696
PGD[40] 0.580 42.181 0.688 0.580 42.298 0.687 0.010 40.543 0.703 0.116 40.567 0.702 0.321 41.397 0.695
StatAT[9] 0.340 33.521 0.634 0.390 28.235 0.629 0.165 39.148 0.602 0.146 45.288 0.535 0.260 36.548 0.600
SimBA[41] 0.475 49.347 0.691 0.830 50.700 0.634 0.010 50.031 0.574 0.161 65.631 0.457 0.369 53.927 0.589
Square[32] 1.000 112.784 0.534 1.000 115.384 0.533 0.605 67.305 0.617 0.623 79.314 0.584 0.807 93.697 0.567
BruSLe[16] 0.925 76.541 0.638 0.820 76.692 0.638 0.605 71.369 0.584 0.347 71.243 0.583 0.674 73.961 0.611
R2BA 1.000 27.352 0.699 1.000 21.066 0.745 0.620 41.380 0.642 1.000 38.681 0.778 0.905 32.120 0.716
TABLE VI: The attack success rate and image quality result on Diffusion-based AIGC detectors
Datasets Attacks FatFormer[11] UnivFD(DRCT)[4] ConvB(DRCT)[4] Average
ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑ ASR↑ BRIS↓ SSIM↑
GenImage[3] FGSM[38] 0.406 53.226 0.643 0.027 53.280 0.626 0.311 53.089 0.622 0.248 53.198 0.630
MIFGSM[39] 0.394 52.903 0.640 0.032 52.930 0.626 0.340 53.005 0.622 0.255 52.946 0.629
PGD[40] 0.406 52.722 0.642 0.043 53.008 0.625 0.311 52.890 0.623 0.254 52.873 0.630
StatAT[9] 0.244 49.824 0.509 0.210 54.794 0.552 0.925 70.335 0.531 0.459 58.318 0.531
SimBA[41] 0.231 57.114 0.522 0.102 45.556 0.334 0.811 46.938 0.497 0.382 49.869 0.451
Square[32] 0.588 83.995 0.523 0.823 105.001 0.519 0.962 107.737 0.489 0.791 98.911 0.510
BruSLe[16] 0.500 70.989 0.536 0.780 69.553 0.553 0.972 70.268 0.548 0.750 70.270 0.546
R2BA 1.000 25.209 0.827 0.898 32.098 0.665 1.000 19.494 0.804 0.966 25.600 0.765
DRCT[4] FGSM[38] 0.302 38.376 0.725 0.020 38.511 0.726 0.005 38.489 0.725 0.109 38.459 0.725
MIFGSM[39] 0.271 38.488 0.726 0.025 38.551 0.724 0.000 38.746 0.724 0.099 38.595 0.725
PGD[40] 0.286 38.158 0.725 0.025 38.384 0.723 0.000 38.103 0.723 0.104 38.215 0.724
StatAT[9] 0.246 35.319 0.547 0.191 40.615 0.609 0.760 45.700 0.592 0.399 40.545 0.582
SimBA[41] 0.196 71.740 0.433 0.010 63.628 0.376 0.565 69.004 0.376 0.257 68.124 0.395
Square[32] 0.618 78.449 0.618 0.628 85.965 0.620 0.875 92.046 0.623 0.707 85.487 0.620
BruSLe[16] 0.523 64.247 0.630 0.568 61.990 0.648 0.490 62.362 0.644 0.527 62.866 0.641
R2BA 1.000 28.764 0.823 0.693 37.029 0.664 0.955 29.609 0.699 0.883 31.801 0.728
Metrics. We evaluate the effectiveness of the attacks using the attack success rate (ASR) and image quality metrics, BRISQUE[42] and SSIM[35]. ASR measures the percentage of AI-generated images misclassified as natural (fake probability<0.5). BRISQUE is a non-reference image quality metric, with higher scores indicating poorer quality, while SSIM measures structural similarity between images, with values closer to 1 indicating better quality.
Basic attacks on white box and black box. We compare four white-box attacks, including FGSM[38], MIFGSM[39], PGD[40] and StatAT[9], and three black-box attacks, including SimBA[41], Square[32] and BruSLe[16]. StatAT[9] is the state-of-the-art (SOTA) white-box post-processing-based attack, and BruSLe[16] is the SOTA black-box query-based attack. For the parameters of baseline attacks, we set the max perturbation magnitude to $20/255$ with the pixel range $[0,1]$, the iterations number to $20$, the max number of queries is $1000$ and use random initialization to start attacks.
Implementation details. About image size, except for attacks on CNNSpot[24] and ResNetND[25] with $256\times 256$ image inputs, all other models use center-cropped $224\times 224$ image inputs to match the suitable input size of the pretrained models. For better analysis, we use bold font to indicate the best result and underline the second best result. We set the max number of queries to 1000 and others are set as follows:
  1. 1.
    Global parameters: $N=100$, $K=10$, $W_{\text{min}}=1$, $W_{\text{max}}=5$, $C_{\text{p}}=C_{\text{total}}=1.5$
  2. 2.
    Gaussian blur: $\alpha\in\{1,3,\dots,13\}$, $\beta\in[0.5,5]$
  3. 3.
    JPEG compression: $\phi\in\{10,\dots,100\}$.
  4. 4.
    Gaussian noise: $\sigma\in[0.0001,(5/255)^{2}]$.
  5. 5.
    Light spot: $\theta\in[0.2,1.8]$, $\gamma\in\{30,\dots,100\}$ and the spot center $(\lambda_{x},\lambda_{y})$ is controlled in the image.
IV-B Anti-detection Performance Analysis
In this subsection, we show the anti-detection performance of attacking AIGC detectors on three datasets. In Table V and Table VI, we present ASR, the image quality metrics BRISQUE (abbreviated as BRIS) and SSIM, and compare the average performance of each metric on all detectors.
Attacks on GAN-based detectors. The experiments in Table V prove that the attack success rate of R2BA reaches the SOTA level, with a max improvement of 37.7% in ASR and an average ASR improvement of about 10%. In terms of image quality, it is at the best level in all the other three detectors except UnivFD[10] and the average results, indicating the excellent invisibility of R2BA. Since we limited the max perturbation magnitude for gradient-based attacks, their image quality on UnivFD is slightly better than ours. However, the attack effectiveness of these white-box methods against the UnivFD detector is nearly zero, confirming that gradient-based white-box attacks struggle with effective anti-detection performance.
Attacks on Diffusion-based detectors. The results in Table VI show that, on GenImage[3] and DRCT[4], R2BA has a maximum improvement of 41.2% in ASR and an average ASR improvement of approximately 18%. Additionally, the image quality metric BRISQUE[42] of R2BA is at an best level, while SSIM[35] is also either close to best or indeed best, demonstrating the excellent invisibility of R2BA in attacking diffusion-based AIGC detectors.
Comprehensive analysis. The results of on three popular AIGC datasets show that perturbations generated by traditional gradient-based or random pixel-modification attacks[16],[38],[39],[40],[41] are often detected as fake by well-trained AIGC detectors. As a post-processing-based white-box attack, due to based on statistical feature consistency, StatAT[9] performs well on facial images but poorly on multi-class natural images, as this type of methods cannot unify the feature dimensions on natural images with inherent multi-class nature and complexity. While Square[32] performs better in ASR but worse in image quality metrics compared to other algorithms, which suggests it may improve the anti-detection performance at the expense of image quality.
In contrast, as a post-processing-based adversarial attack, R2BA does not require consideration of feature-level consistency. Instead, it targets the vulnerabilities of the detectors to achieve efficient attacks, making it an adversarial attack method that is independent of the image classes. The results indicate that, by optimizing the post-processing fusion intensity, R2BA achieves impressive and comprehensive anti-detection performance. It successfully evades both GAN-based and diffusion-based AIGC detectors. The realistic-like adversarial examples generated by R2BA, matching the pattern of the real-world images, lead the detectors to recognize them as real natural images. In other words, when dealing with post-processing fusion attacks, the existing AIGC detectors have significant performance degradation, which suggests that if the AIGC detectors are put into practical application, their security cannot be guaranteed and remains to be further explored.
IV-C Robustness Comparison Experiments
The robustness of adversarial examples to common image post-processing techniques is crucial for determining their real-world applicability. If an adversarial example is vulnerable to such techniques, its adversariality is erased, and the threat to AIGC detectors is eliminated.
TABLE VII: Robustness experiments against known post-processing — JPEG compression and Gaussian noise
Dataset Attacks ASR↑ - JPEG compression ASR↑ - Gaussian noise
50 60 70 80 90 AVG 1/255 2/255 3/255 4/255 5/255 AVG
CNNSpot[24] FGSM[38] 0.142 0.190 0.218 0.268 0.262 0.216 0.325 0.330 0.334 0.345 0.357 0.338
MIFGSM[39] 0.152 0.185 0.207 0.269 0.260 0.214 0.318 0.320 0.336 0.336 0.358 0.334
PGD[40] 0.139 0.187 0.215 0.263 0.256 0.212 0.329 0.328 0.331 0.335 0.350 0.335
StatAT[9] 0.263 0.322 0.224 0.228 0.219 0.251 0.206 0.217 0.240 0.297 0.326 0.257
SimBA[41] 0.149 0.213 0.259 0.324 0.361 0.261 0.314 0.255 0.205 0.189 0.186 0.230
Square[32] 0.521 0.527 0.538 0.568 0.640 0.559 0.594 0.553 0.543 0.540 0.525 0.551
BruSLe[16] 0.204 0.233 0.279 0.355 0.437 0.301 0.500 0.431 0.394 0.390 0.381 0.419
R2BA 0.720 0.753 0.801 0.837 0.875 0.797 0.706 0.598 0.575 0.553 0.547 0.596
GenImage[3] FGSM[38] 0.237 0.233 0.241 0.321 0.427 0.292 0.242 0.195 0.200 0.167 0.139 0.189
MIFGSM[39] 0.182 0.303 0.218 0.323 0.419 0.289 0.242 0.206 0.180 0.164 0.118 0.182
PGD[40] 0.320 0.290 0.282 0.276 0.444 0.322 0.253 0.228 0.181 0.171 0.154 0.197
StatAT[9] 0.554 0.585 0.366 0.375 0.476 0.471 0.394 0.392 0.371 0.375 0.382 0.383
SimBA[41] 0.454 0.323 0.350 0.293 0.372 0.359 0.295 0.249 0.244 0.205 0.208 0.240
Square[32] 0.493 0.478 0.387 0.420 0.542 0.464 0.605 0.555 0.519 0.459 0.436 0.515
BruSLe[16] 0.385 0.480 0.408 0.536 0.485 0.459 0.592 0.498 0.437 0.346 0.284 0.431
R2BA 0.654 0.691 0.699 0.748 0.793 0.717 0.712 0.577 0.536 0.474 0.402 0.540
DRCT[4] FGSM[38] 0.112 0.150 0.117 0.188 0.188 0.151 0.104 0.103 0.094 0.075 0.075 0.090
MIFGSM[39] 0.139 0.142 0.104 0.155 0.191 0.146 0.099 0.096 0.087 0.093 0.075 0.090
PGD[40] 0.090 0.125 0.115 0.177 0.181 0.138 0.099 0.096 0.078 0.081 0.074 0.085
StatAT[9] 0.374 0.386 0.327 0.400 0.455 0.388 0.344 0.328 0.320 0.323 0.297 0.322
SimBA[41] 0.353 0.352 0.302 0.177 0.338 0.304 0.204 0.191 0.198 0.188 0.177 0.192
Square[32] 0.374 0.344 0.337 0.177 0.423 0.331 0.447 0.385 0.349 0.331 0.326 0.368
BruSLe[16] 0.201 0.219 0.189 0.261 0.261 0.226 0.337 0.256 0.210 0.177 0.175 0.231
R2BA 0.578 0.594 0.628 0.697 0.740 0.647 0.661 0.555 0.508 0.478 0.432 0.527
TABLE VIII: Robustness experiments against unknown post-processing — Rotation and Resize
Dataset Attacks ASR↑ - Rotation ASR↑ - Resize
2 4 6 8 10 AVG 1/2 3/4 5/4 3/2 7/4 AVG
CNNSpot[24] FGSM[38] 0.356 0.386 0.396 0.406 0.432 0.395 0.145 0.247 0.248 0.248 0.248 0.227
MIFGSM[39] 0.365 0.389 0.407 0.415 0.430 0.401 0.149 0.240 0.254 0.244 0.250 0.227
PGD[40] 0.362 0.387 0.395 0.412 0.429 0.397 0.147 0.243 0.256 0.255 0.256 0.231
StatAT[9] 0.257 0.279 0.315 0.362 0.396 0.322 0.248 0.232 0.225 0.225 0.230 0.232
SimBA[41] 0.325 0.336 0.345 0.353 0.366 0.345 0.021 0.080 0.133 0.131 0.136 0.100
Square[32] 0.494 0.494 0.500 0.500 0.503 0.498 0.537 0.550 0.582 0.576 0.587 0.566
BruSLe[16] 0.331 0.346 0.372 0.386 0.396 0.366 0.104 0.250 0.338 0.343 0.345 0.276
R2BA 0.761 0.719 0.705 0.690 0.533 0.681 0.771 0.827 0.844 0.833 0.832 0.821
GenImage[3] FGSM[38] 0.008 0.022 0.049 0.052 0.042 0.035 0.178 0.231 0.258 0.278 0.265 0.242
MIFGSM[39] 0.012 0.027 0.037 0.045 0.051 0.034 0.165 0.228 0.247 0.254 0.256 0.230
PGD[40] 0.013 0.022 0.035 0.059 0.049 0.036 0.159 0.228 0.235 0.256 0.241 0.224
StatAT[9] 0.218 0.225 0.210 0.215 0.159 0.206 0.118 0.276 0.233 0.220 0.219 0.213
SimBA[41] 0.120 0.168 0.200 0.210 0.220 0.184 0.089 0.146 0.107 0.104 0.100 0.109
Square[32] 0.245 0.302 0.310 0.322 0.335 0.303 0.310 0.319 0.308 0.307 0.308 0.310
BruSLe[16] 0.015 0.015 0.022 0.027 0.030 0.022 0.096 0.182 0.103 0.108 0.096 0.117
R2BA 0.544 0.518 0.513 0.500 0.428 0.500 0.439 0.613 0.684 0.679 0.678 0.619
DRCT[4] FGSM[38] 0.126 0.190 0.191 0.179 0.194 0.176 0.265 0.379 0.258 0.278 0.265 0.289
MIFGSM[39] 0.122 0.188 0.208 0.205 0.176 0.180 0.264 0.366 0.247 0.254 0.256 0.277
PGD[40] 0.126 0.182 0.191 0.166 0.172 0.167 0.280 0.358 0.235 0.256 0.241 0.274
StatAT[9] 0.332 0.298 0.297 0.303 0.286 0.303 0.239 0.368 0.338 0.329 0.343 0.323
SimBA[41] 0.173 0.170 0.194 0.222 0.226 0.197 0.158 0.233 0.193 0.234 0.208 0.205
Square[32] 0.410 0.413 0.410 0.432 0.423 0.417 0.460 0.497 0.493 0.521 0.499 0.494
BruSLe[16] 0.181 0.168 0.192 0.191 0.166 0.180 0.263 0.367 0.285 0.291 0.281 0.297
R2BA 0.622 0.614 0.551 0.531 0.435 0.551 0.495 0.703 0.753 0.753 0.750 0.691
Therefore, we further conduct robustness comparison experiments for these adversarial attacks. Since our method is designed based on four post-processing methods, it is naturally robust to common platform image post-processing. Table VII shows robustness experiments against known post-processing, JPEG compression and Gaussian noise, while Table VIII compares robustness against unknown post-processing, Rotation and Resize (not included in the four post-processing we selected), offering a more comprehensive analysis. AVG in both tables refers to average performance.
Against known image post-processing. Apparently, The experimental results in Table VII show that R2BA far outperforms other adversarial attacks at most settings in robustness. This result indicates that the adversarial examples generated by R2BA can better maintain their adversariality in practical applications and have stronger realistic adaptability. This results also suggest the effectiveness of our post-processing fusion optimization method. Due to the post-processing fusion optimization used in the attack, the image with the best post-processing fusion intensity generated by R2BA, when subjected to another similar post-processing, will deviate from the best post-processing intensity, which theoretically leads to a certain degree of performance degradation in terms of anti-detection. And so it is. The robustness experiments in Table VII further confirm this, providing additional evidence for the effectiveness of optimization in R2BA.
Against unknown image post-processing. In Table VIII, it can be seen that our method remains remarkably robust when dealing with the unknown post-processing. This is because, essentially, our proposed method, R2BA, is not a simple additive noise perturbation and will not have its adversariality erased by simple image post-processing techniques, such as rotation or resize. In contrast, apart from causing severe image distortion[32], adversarial attack methods based on simple additive noise perturbations are easily faded or eliminated after image post-processing, leading to attack failure.
Overall, after image processing, although the anti-detection performance is some degraded, the ASR of R2BA remains around $0.5$ or even higher. And when a detector’s detection success rate is only 50%, its detection is unreliable. On the other hand, the robustness experiments demonstrate that stronger post-processing intensity is not always better. This is consistent with our discovery from the pre-experiments in Fig. 2 of subsection III-A. The adversarial examples generated with excessively strong post-processing intensity are heavily distorted, which in turn causes the detector to classify the post-processed AIGC image as fake. This also proves that, to some extent, the robustness of AIGC detectors to widely-used image post-processing techniques is uncertain and fluctuates, indicating a potential security risk.
IV-D Visual Comparison and Analysis
Figure 4: Comparative experiments on image quality of adversarial examples. The dataset we use is GenImage[3] and the detector is FatFormer[11]. We zoom in on the local details at the end and visualize the perturbations across the entire image.
The invisibility of the adversarial example is crucial in determining whether this perturbation is likely to be erased. Better visualisation increases the likelihood of this adversarial example being shared in real life or on the internet. Therefore, we provide a visual comparison in Fig. 4. Ten images from different classes (e.g., dogs, landscapes, cars, letters, etc.) are randomly chosen for the image quality comparison, and one additional image is selected for detailed comparison and perturbation visualization. Combined with the excellent anti-detection performance of R2BA analysed in subsection IV-B, we can observe that R2BA also maintains quite good image quality, demonstrates strong image realism, and achieves multi-class anti-detection performance on natural images.
It is worth noting that, to ensure objectivity and fairness, some images with obvious post-processing traces (e.g., visible light spot) are specially selected for comparison. From both the detailed comparison and perturbation visualization, our image quality is clearly superior. Moreover, the post-processed AIGC images have practical significance. For example, our light spot can simulate real-world flashlight, while JPEG compression and Gaussian blur mimic real-world image quality degradation. These modifications enhance the realism of adversarial examples, whether viewed from a human or a detector’s perspective. Additionally, the visualisation also shows that existing AIGC detectors are not robust in detecting post-processed AIGC images, and slight image post-processing operations may cause the detector to output incorrect results.
IV-E Black-box Query Count Comparison Experiments
TABLE IX: Black box Average queries on FatFormer
Detector Attack CNNSpot[24] GenImage[3] DRCT[4]
FatFormer[11] SimBA[41] 971.055 886.475 960.372
Square[32] 178.402 339.875 213.075
BruSLe[16] 791.555 555.690 657.015
R2BA 166.834 119.139 132.724
In practice, the number of queries to the detector is limited. A smaller number of queries can save time and control the budget. Therefore, as a black-box query-based attack, the number of queries is an important factor to consider. To ensure fairness, we check whether the adversarial point is reached after every 100 queries (we use 100 particles, where 1 particle represents 1 query). In Table IX, we use FatFormer[11] as the detector, which is the only one that covers all three datasets and generative models in its research and maintains good detection results for all 3 datasets. Evidently, the number of queries for R2BA is significantly smaller than for other black-box attacks. Therefore, compared to other query-based adversarial attacks, R2BA is able to maintain excellent anti-detection performance with fewer queries.
IV-F Effect of Particle Number
To further explore the influence of particle number on R2BA optimization, we vary the number of particles and the number of iterations to compare the variations in in ASR, BRISQUE[42], SSIM[35], and the number of queries under different particle numbers. To ensure fairness of the experiments, the maximum number of queries is always kept at 1000.
The experiments are shown in Fig 5. We observe that when the number of particles increases, the ASR also increases. When the number of particles reaches 100, the ASR of all detectors reaches a quite high level. However, the effect of the particle number on BRISQUE and SSIM is not significant. This indicates that when the intensity of post-processing is controlled within a certain range, the image quality of the adversarial examples generated by R2BA does not fluctuate significantly with changes in other parameters of the algorithm.
Interestingly, in the experiment of UnivFD(DRCT), although we typically believe that an increase in the number of particles leads to a rise in the number of queries, the experimental results show that the number of queries is actually higher with a lower number of particles. Building on the previous experimental results regarding attack success rate, we know that UnivFD(DRCT) is a detector that is harder to attack. This suggests that when attacking robust detectors, our optimization algorithm struggles to find a suitable optimization direction with a lower number of particles, and thus more queries are needed for exploration. On the other hand, when the number of particles is higher, the query cost for a single iteration increases, and the increase in the query base leads to a multiplicative trend in the number of queries.
Therefore, choosing the right number of particles is crucial for optimizing the attack success rate of the algorithm. From the experimental results, it can be seen that choosing 100 particles and carrying out 10 iterations of the attack is a more ideal choice, as it achieves a better balance between the attack success rate and the number of queries.
Figure 5: The effect of the number of particles on the attack success rate, BRISQUE, SSIM, and the number of queries.
IV-G Ablation Experiments
TABLE X: Ablation experiments on the optimization algorithm and post-processing techniques
Dataset Attacks FatFormer[11] UnivFD(DRCT)[4] ConvB(DRCT)[4] Average
ASR↑ BRIS↓ SSIM↑ Query↓ ASR↑ BRIS↓ SSIM↑ Query↓ ASR↑ BRIS↓ SSIM↑ Query↓ ASR↑ BRIS↓ SSIM↑ Query↓
GenImage[3] Random 0.244 33.186 0.711 N/A 0.005 31.284 0.726 N/A 0.292 31.915 0.725 N/A 0.181 32.128 0.721 N/A
w GB 0.781 68.481 0.708 322.5 0.789 73.279 0.699 298.4 0.704 51.660 0.767 453.7 0.758 64.473 0.725 358.2
w JPEG 1.000 36.360 0.878 131.2 0.676 34.715 0.905 401.1 0.787 36.341 0.878 376.9 0.821 35.805 0.887 303.0
w GN 0.538 24.743 0.883 576.9 0.249 23.233 0.884 805.4 0.602 22.807 0.868 564.8 0.463 23.594 0.878 649.0
w LS 0.144 24.394 0.959 898.8 0.141 23.469 0.965 886.5 0.315 24.193 0.943 813.9 0.200 24.019 0.956 866.4
w/o GB 1.000 21.328 0.835 175.6 0.694 35.110 0.771 442.5 0.934 27.846 0.759 255.7 0.876 28.095 0.789 291.3
w/o JPEG 0.944 23.841 0.693 203.1 0.876 30.768 0.643 272.6 0.962 50.374 0.706 234.9 0.927 34.994 0.681 236.9
w/o GN 1.000 40.689 0.780 135.0 0.882 44.357 0.701 272.8 0.991 58.827 0.747 147.2 0.957 47.958 0.743 185.0
w/o LS 1.000 24.193 0.724 148.8 0.806 30.665 0.624 344.1 0.972 33.270 0.714 234.0 0.926 29.376 0.687 242.3
R2BA 1.000 25.209 0.827 129.1 0.898 32.098 0.665 261.8 1.000 19.494 0.804 112.3 0.966 25.600 0.765 167.7
To comprehensively validate the effectiveness of R2BA, we conducted ablation experiments, as shown in Table X. The experiments are divided into three main parts:
Randomly generated post-processing intensity. We use randomly generated post-processing intensity (without optimization, so the query is N/A) to create adversarial examples for validating the effectiveness of R2BA optimization. The results show that without guidance from the detector’s fake probability, these examples struggle to cross decision boundaries and perform poorly in anti-detection. If the randomly generated intensity is weaker, the adversariality is also weaker, while stronger intensity will lead to severe distortion, prompting the detector to classify the image as AI-generated as well.
R2BA with single post-processing. Due to the different vulnerabilities of detectors to various post-processing techniques, there are performance differences between R2BA with different post-processing. For example, the ASR of R2BA with Gaussian noise is 0.249 on UnivFD(DRCT)[4], but nearly 3 times higher on ConvB(DRCT)[4]. In addition, the anti-detection performance of R2BA with JPEG compression significantly better than that of R2BA with light spot, likely because detectors are more sensitive to high-frequency information (which JPEG compression causes to be lost) than to changes in statistical features (which light spot alters).
R2BA without certain post-processing. The anti-detection performance of R2BA without certain post-processing slightly decreases. This is because the effects of post-processing on features are not mutually exclusive. For example, JPEG compression leads to the loss of high-frequency information but also indirectly affects statistical features. Although optimization becomes harder, R2BA without certain post-processing can still generate effective adversarial examples to achieve successful attacks, at the cost of more queries. As a result, R2BA without certain post-processing shows lower anti-detection performance and requires more queries compared to R2BA with all four post-processing.
Considering all metrics, and comparing with ablated adversarial attack methods, R2BA with 4 post-processing fusion attacks achieves the best results in both attack success rate and query count, while the image quality metrics BRISQUE[42] and SSIM[35] are either the best or close to the best.
IV-H Attacks on Commercial AIGC Detectors
TABLE XI: Attacks on Alibaba’s Commercial AIGC Detection
Attacks CNNSpot[24] GenImage[3] DRCT[4] Average
SR↓ BRIS↓ SSIM↑ Query↓ SR↓ BRIS↓ SSIM↑ Query↓ SR↓ BRIS↓ SSIM↑ Query↓ SR↓ BRIS↓ SSIM↑ Query↓
Clean 1.000 N/A N/A N/A 1.000 N/A N/A N/A 1.000 N/A N/A N/A 1.000 N/A N/A N/A
FGSM[38] 1.000 40.071 0.704 N/A 0.980 53.280 0.626 N/A 1.000 38.511 0.726 N/A 0.993 43.954 0.685 N/A
MIFGSM[39] 1.000 40.624 0.703 N/A 0.970 52.930 0.626 N/A 1.000 38.551 0.724 N/A 0.990 44.035 0.685 N/A
PGD[40] 1.000 40.543 0.703 N/A 0.970 53.008 0.625 N/A 1.000 38.384 0.723 N/A 0.990 43.978 0.684 N/A
StatAT[9] 0.230 39.148 0.602 N/A 0.165 54.794 0.552 N/A 0.135 40.615 0.609 N/A 0.177 44.852 0.587 N/A
SimBA[41] 0.295 69.677 0.724 97.9 0.745 64.603 0.666 98.9 0.750 66.595 0.716 99.9 0.597 66.958 0.702 98.9
Square[32] 0.000 113.216 0.535 20.0 0.000 137.065 0.457 20.0 0.005 112.051 0.562 20.4 0.002 120.777 0.518 20.1
BruSLe[16] 0.185 90.584 0.578 45.9 0.505 86.623 0.486 68.7 0.440 77.847 0.569 66.8 0.377 85.018 0.544 60.5
R2BA 0.000 25.593 0.740 20.0 0.000 25.410 0.691 20.0 0.000 29.640 0.733 20.0 0.000 26.881 0.722 20.0
We use Alibaba’s commercial AIGC detection API[43] to compare the effectiveness of the attacks under real-world conditions. For budget and fairness, we set the maximum query to 100, the number of particles in R2BA to 10, and check every 20 queries to determine if the attack succeeds. For white-box attacks, we test adversarial examples generated by attacking the UnivFD[10] and UnivFD(DRCT)[4] detectors, the two most robust detectors across three datasets.
In Table XI, clean refers to the detection success rate (SR) of clean AIGC images, which is 100% for all cases. N/A in clean means irrelevant, and in white-box, N/A means query not applicable. Furthermore, in Fig. 6, we present a visual comparison of two high-performance attacks[9],[32]. Overall, the experiments show that RBA exhibits excellent anti-detection performance, high invisibility, and outstanding query efficiency in real-world scenarios. The experiment further demonstrates that AIGC detectors can classify post-processed AIGC images as User-generated images.
Figure 6: Visual comparison on Alibaba’s commercial AIGC detection. The red box indicates AIGC and the green box indicates UGC. The labels for AIGC and UGC come from Alibaba AIGC detection API[43].
V Conclusion
In this paper, we explore the robustness of GAN-based and diffusion-based AIGC detectors under adversarial example attacks. Given the uncertainty of detectors in real-world scenarios and the discovery that detectors vary in vulnerability to different types of post-processing, we propose a realistic-like robust black-box adversarial attack method, R2BA, based on post-processing fusion optimization. Our method generates adversarial examples that are both robust and maintain high image quality by optimizing post-processing intensity, effectively evading both GAN-based and diffusion-based detection. We conduct extensive experiments on three GAN-based and diffusion-based AIGC datasets, as well as four GAN-based AIGC detectors, three diffusion-based AIGC detectors and one commercial AIGC detection API. The comprehensive experimental results demonstrate that the proposed attack method achieves promising anti-performance in both GAN-based and diffusion-based settings, under the original scenario and the robust scenario, respectively.